x2 y2 satisfies the differential equation Show that z = (x2+y2) az az y X дх 2z а. ay

please give me whole solution for this question
Answer #1

xyz +11) 27y2 g(x) f(x) = f(2). (902)) = g(8). g(x) 92-2?y? 2 (2+y?) -(a+ %). 2. (2+2) say? (22) - (2²+} ?(21) - 22?y? - 22LHS LHS 222 + y 22 00 160223g=y) + (p*y 2-97a13 = 22y? 236 – 2.354+2744 24%24_24922 = 22\y?= 241-2 = 2 (à ey - 2 | 1 ( 4 ))​​​​​​